OpenStudy (anonymous):

Find the following derivative ln(x)[e^((sin^2)(x))]

OpenStudy (anonymous):

(e^(sin^2(x)) (1+x log(x) sin(2 x)))/x Use the product rule. lnx =1/x and e^... = Derivative of ... * original e^... equation.